# The Best Instruction-Tuning Data are Those That Fit

### v2

@@ Abstract
High-quality supervised finetuning (SFT) data are crucial for eliciting strong capabilities from pretrained large language models (LLMs). Typically, instructions are paired with multiple responses sampled from other LLMs, which are often out of the distribution of the target model to be finetuned. This, at scale, can lead to diminishing returns and even hurt the models’ performance and robustness. We propose GRAPE, a novel SFT framework that accounts for the unique characteristics of the target model. For each instruction, it gathers responses from various LLMs, and selects the one with the highest probability measured by the target model, indicating that it aligns most closely to the target model’s pretrained distribution; it then proceeds with standard SFT training. We first evaluate GRAPE with a controlled experiment, where we sample various solutions for each question in UltraInteract from multiple models and finetune commonly used LMs like Llama3.1-8B, Mistral-7B and Qwen2.5-7B on GRAPE-selected data. GRAPE significantly outperforms strong baselines, including distilling from the strongest model with absolute gain up to 13.8% averaging across benchmarks, and training on 3$\times$ more data with maximum 17.3% performance improvements. GRAPE’s strong performance generalizes to realistic settings. We experiment with the post-training data used for Tulu3 and Olmo-2. GRAPE can outperform strong baselines with 4.5 times the data by 6.1% and state-of-the-art data selection approach by 3.9% on average performance. Remarkably, using 1/3 data and half number of epochs, GRAPE allows Llama3.1-8B to surpass the performance of Tulu3-SFT by 3.5%.
Machine Learning, ICML
@@ 1 Introduction
High-quality, large-scale supervised data is crucial for supervised instruction finetuning (SFT; Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024a; Zheng et al., 2024). A common practice of data collection involves sampling responses from strong language models, predominantly focusing on expanding the size of the dataset and improving the overall quality of the responses (Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024c; Chen et al., 2024). However, recent research suggests that there is more complex dynamics involved (Xu et al., 2024d). A plateau effect in synthetic data scaling, where performance either stagnates or even declines as the size of the synthetic data increases beyond a certain point, has been widely observed. This phenomenon arises due to issues such as diminishing diversity (Padmakumar & He, 2024; Guo et al., 2023) and distortion in the data distribution (LeBrun et al., 2021), which ultimately undermine the base model’s performance and robustness (Alemohammad et al., 2024; Gerstgrasser et al., 2024; Shumailov et al., 2023; Dohmatob et al., 2024; Hataya et al., 2023; Martínez et al., 2023a, b; Bohacek & Farid, 2023; Briesch et al., 2023).
@@
Thus, effective instruction tuning requires more than scaling up the data; it often needs “tailoring” the data to the unique characteristics of the target model. Existing works focus on enhancing the model’s existing knowledge and capabilities (Du et al., 2023) and optimizing the curriculum progression for instruction tuning (Zhao et al., 2024b; Lee et al., 2024a; Feng et al., 2023; Setlur et al., 2024). They typically prescribe questions for each model to learn solutions. However, it remains elusive which responses best suit a given base model for fine-tuning. Meanwhile, tailoring the responses to the model has been a crucial ingredient for the success of later phases of LLM development, particularly through on-policy preference learning (Tajwar et al., 2024; Zhang et al., 2024c, a; Miao et al., 2024; Gulcehre et al., 2023; Azar et al., 2023a; Tang et al., 2024a; Zhuang et al., 2023), and on-policy/online reinforcement learning (Guo et al., 2024b; Liu et al., 2024d; Zhou et al., 2024c).
@@
Inspired by these insights, we hypothesize that SFT can similarly benefit from aligning data with the model, the core idea behind GRAPE. For each instruction, GRAPE gathers and selects response(s) from various sources that are closest to the target model’s pretrained distribution. This is achieved by calculating the probability of each response using the target model and selects the one with the highest probability(§3).After obtaining these in-distribution responses, one can proceed with standard SFT without any modification to the training. Unlike existing datasets that usually contains one-size-fits-all responses for each instruction without customization (Yu et al., 2024; Yuan et al., 2024; Lian et al., 2023b; Teknium, 2023), GRAPE curates model-dependent SFT datasets that better matches the base model’s distribution, better mitigating the risks associated with distribution shift like spurious correlations and catastrophic forgetting, while posing minimum overhead.
@@
To GRAPE’s advantage, many existing datasets share overlapping instructions but contain different high-quality responses, e.g., the instructions in Flan (Longpre et al., 2023), GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and the post-training recipes that re-use SFT instructions for preference learning (Lambert et al., 2024; OLMo et al., 2025). Therefore, GRAPE can directly select, for each model, the best fit(s) among the off-the-shelf responses without producing new responses, which we do in the experiments(§ 5).
We first validate our hypothesis through extensive controlled experiments on a reasoning dataset - UltraInteract-SFT (Yuan et al., 2024) and demonstrate the importance of supervising base models with in-distribution responses(§4). We experimented on 4 popular pretrained LMs from Mistral (MistralAI, 2024c), Llama3.1 (Dubey et al., 2024) and Qwen2.5 (Hui et al., 2024) families. Notably, models fine-tuned with GRAPE-selected responses outperform those trained on a 3$\times$ larger datasets up to $17.3\%$ absolute gain on average performances, and responses from the strongest model under consideration - Llama3.1-405B-Instruct by significant levels. We then experiment in realistic scenarios, collecting responses from post-training data for Tulu3 and Olmo-v2. We demonstrate the effectiveness of GRAPE by outperforming training on all these available data that is 4.5 times the size of data curated by GRAPE by up to 6.1%. Remarkably, GRAPE allows finetuning a Llama3.1-8B base model to exceed the performance of Tulu-8B-SFT using 1/3 data and half number of epochs.
Our results prove GRAPE as a simple but effective and scalable approach to improving supervised fine-tuning by aligning training distribution and that of the pre-trained checkpoint.
@@ 2 Background and Motivation
The investigation of this work into the distribution match between the pre-trained LM and supervised fine-tuning (SFT) data is inspired by recent findings on policy optimization for LM alignment with RL (Ouyang et al., 2022) and preference learning (Rafailov et al., 2023; Ethayarajh et al., 2024). While the importance of matching training data distribution with the policy has been well noted in both traditional RL (Shi et al., 2023; Fujimoto et al., 2018; Kumar et al., 2019; Peng et al., 2019; Wang et al., 2021; Arora & Goyal, 2023; Jiang & Li, 2016; Tang & Abbeel, 2010) and LM settings (Xiong et al., 2024), preference learning algorithms like DPO (Rafailov et al., 2023), IPO (Azar et al., 2023b) and KTO (Ethayarajh et al., 2024) first emerged as off-policy algorithms. However, subsequent research has highlighted the performance gap between on-policy and off-policy training due to distribution shifts  (Xu et al., 2024b; Tang et al., 2024b) and proposed various mitigation strategies  (Zhuang et al., 2023; Zhou et al., 2024c; Zhang et al., 2024a; Xiong et al., 2024; Guo et al., 2024b), showing that training models on data more closely aligned with their policy distribution can significantly improve performance, while failing to do so can yield sub-optimal policies or those that are harder to generalize.
@@
The base distribution of pre-trained language models—shaped by extensive training on vast and diverse datasets—is inherently robust and generalizable. Therefore, during supervised fine-tuning phase, the pre-trained distribution should be carefully preserved (Kumar et al., 2022; Cohen-Wang et al., 2024; He et al., 2023; Yang et al., 2024d; Ding et al., 2023), to best retain the knowledge and capabilities that emerge during pre-training (Zhou et al., 2023). If the proximity between the pre-trained distribution and the fine-tuning data is not maintained, the limited number of training examples available during SFT, compared to the vast scale of pre-training data, can increase the risk of distribution distortion. This misalignment can lead to issues such as catastrophic forgetting (Aghajanyan et al., 2020; Yang et al., 2024d) and the emergence of spurious correlations (Feldman, 2021).
@@
We build on principles of on-policy alignment techniques with key distinctions tailored for SFT. Given that SFT reprsents a more primitive stage of post-training than preference learning, strictly sampling responses from the base model itself can introduce risks such as instability, bias reinforcement, knowledge stagnation, and overfitting (Herel & Mikolov, 2024; Mobahi et al., 2020; Allen-Zhu & Li, 2023). To address this, we advocate for an approach that stays in-distribution while delivering effective supervision to the base model. To this end, we propose to gather and select responses from various sources, and select one that is closest to the target model’s pretrained distribution, which we name GRAPE.
@@ 3 Methodology
We introduce GRAPE, a surprisingly simple yet effective methodology to enhance supervised fine-tuning (SFT) by customizing the training data for the base model. The key idea is to find responses among a candidate pool for each instruction $x_{i}$ such that they align closely with the base model’s pretrained distribution $\pi_{\theta_{0}}$.
@@
  1. 1.
    Response Collection (§3.1) Collect a pool of high-quality candidate responses either from existing datasets or sampling from multiple LLMs.
  2. 2.
@@ 3.1 Collecting Responses from Existing Resources
For instruction-tuning of language models, high-quality instructions are more difficult to collect than responses (Xu et al., 2024c; Liu et al., 2024a). Therefore, it is a common practice to reuse existing instruction-tuning prompts while generating diverse responses using various methods tailored to specific requirements. For instance, instructions from Flan (Longpre et al., 2023), OpenOrca (Lian et al., 2023a), ShareGPT (Team, 2023), and the training splits of GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and CodeContests (Li et al., 2022) are frequently reused in datasets like Olmo (OLMo et al., 2025), Tulu (Lambert et al., 2024), OpenHermes (Teknium, 2023), OpenOrca (Lian et al., 2023a), MetaMath (Yu et al., 2024), MathInstruct (Yue et al., 2023), UltraFeedback (Cui et al., 2024), and UltraInteract (Yuan et al., 2024), whether for SFT or preference learning. The solutions are generated using different models or follow varying styles depending on the specific needs. This naturally leads to a situation where a single instruction with multiple responses becomes a readily available resource. GRAPE therefore leverages these pre-existing response candidates to tailor training dataset that better aligns with the base model’s distribution. When such resources are unavailable or insufficient, practitioners can generate new responses and apply GRAPE.
Responses for the $i$-th instruction are collected from various datasets to form $\mathcal{R}_{i}=\{y{{}_{i}^{j}}:j=1,\dots,J^{(i)}\}$, where $J^{(i)}$ denotes the number of responses collected for the $i$-th instruction.
@@ 3.2 Customize Dataset For Models
We then compute the conditional probability of each response $\pi_{\theta_{0}}(y{{}_{i}^{j}}\mid x_{i})$. We rank the responses based on the conditional probability and take those with the highest probability for each instruction.
Notably, GRAPE’s selection process requires only a forward pass through the candidates and does not require gradient computation. Its overhead is substantially lower than many model-based data selection algorithms (Xia et al., 2024; Yang et al., 2024b; Liu et al., 2024b; Zhao et al., 2021; Zhang et al., 2024b; Pan et al., 2024).
Additionally, it is important to distinguish GRAPE from the other preplexity-based data selection and curriculum planning methods (Wu et al., 2024; Li et al., 2024b; Liu et al., 2024c). Existing approaches focus on selecting instructions by using perplexity as a difficulty measure, which differs from GRAPE that uses probability to select for each instruction in a fixed instruction set, responses that better matches with the base model’s distribution. Our experiments in §5 demonstrate that low-probability responses with fixed set of instructions are detrimental to performance, further emphasizing the fundamental difference in the two processes.
4 Preliminary Experiments
In this section, we seek to empirically verify the hypothesis in §2. To evaluate it under a controlled set-up, we consider using instructions from UltraInteract-SFT, and constrain within chain-of-thought reasoning (predominantly for coding, logic and math.) scenario. This allows us to control confounding factors like style and format while ensuring manageable data quality control during data collection. The insights from this experiment will inform the later more realistic setting in §5.
@@ 4.1 Experimental Setup
We detail our experiment settings below.
4.1.1 Training Data Curation
In this controlled experiment, we focus on chain-of-thought reasoning (Wei et al., 2022; Wang et al., 2024; Luo et al., 2024; Cobbe et al., 2021; Li et al., 2023; Lightman et al., 2023). Different models may follow different reasoning paths to solve a problem, while their final solutions can be easily verified.
Figure 2: The training loss curve of Llama3.1-8B on Tulu-3-Olmo-2 combo. Best stands for training with highest-probability responses (GRAPE-selected); Random stands for randomly chosen responses for each instruction in the pool, Worst stands for lowest probability responses in the base model’s distribution. Throughout training, Best has lower loss than Random than Worst.
We use UltraInteract-SFT (Yuan et al., 2024), which contains approximately $80,800$ unique instructions covering coding, math (chain-of-thought and program-aided) and logic reasoning domains , where each instruction is paired with varying numbers of different responses to contain a total $>280,000$ training examples. The responses in the dataset are strictly in step-wise format. For each instruction, we use GRAPE to select the top-ranking responses, ensuring that the number of responses matches the original UltraInteract-SFT dataset for fair comparisons.
Response Generation
We collect responses from a diverse set of models of various sizes across model families, including Mixtral-7x7B-Instruct (Jiang et al., 2024), Codestral-22B (MistralAI, 2024a), Mistral-Small (MistralAI, 2024b), Llama-3.1-70B-Instruct and Llama-3.1-405B-Instruct (Dubey et al., 2024), and Qwen2.5-72B-Chat (Yang et al., 2024a), resulting in approximately 10x additional responses per instruction. The responses are then filtered based on the answers to ensure their validity following Yuan et al. (2024).
Figure 3: The average performance curve of Llama3.1-8B on Tulu-3-Olmo-2 combo. Best stands for training with highest-likelihood responses (GRAPE-selected); Random stands for randomly chosen responses for each instruction in the pool. The average is taken over all items on table 2 except AlpacaEval2.
4.1.2 Base Models
To demonstrate the generalizability of GRAPE, we evaluate its performance across multiple LLMs, including Llama-3.1-8B and Llama-3.2-3B from Llama-3 (Grattafiori et al., 2024) family, Mistral-7B (Jiang et al., 2023) and Qwen2.5-7B (Hui et al., 2024).
4.1.3 Evaluation
@@
4.1.4 Baselines
We compare GRAPE with several baselines. Except for Up-scaled Dataset which is 3×3$\times$3 × larger, the rest are controlled for the number of responses per-instruction to be the same as UltraInteract-SFT.
  • Original dataset. Performing SFT over the original UltraInteract-SFT dataset as is. The original dataset also contains high-quality, verified responses.
    We compare GRAPE to directly using a standard, well-curated dataset. Responses from a powerful model are likely to represent an upper-bound quality of training data, and comparing against this baseline allows us to isolate whether further improvements stem from the strategic response selection performed by GRAPE.
  • Responses from the strongest model under consideration. We use responses exclusively from the best model used to generate responses (in our case, Llama3.1-405B-Instruct).
  • Up-scaled Dataset. The comparison with a larger dataset assesses how well selection improves instruction tuning compared to training on a larger set of diverse correct responses. Each instruction is paired with three times of randomly selected responses from the pool. [^1] For example, if UltraInteract contains 3 different responses for instruction $x$, we end up including 9 for this setting.
We train all models for 1 epoch with a learning rate of $10^{-5}$.
@@ 4.1.4 Baselines
Qwen2.5-7B Ours 77.4 48.9 70.7 56.4 45.3 67.7 66.3 37.4 40.1 56.7 -
Table 1: Result of synthetic experiment on UltraInteract-SFT. The last column, Abs. $\Delta$ is the absolute improvement of GRAPE over that row.

### v3

@@ Abstract
[…]
@@ 1 Introduction
High-quality, large-scale supervised data is crucial for supervised fine-tuing(SFT; Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024a; Zheng et al., 2024). A common practice of collecting SFT data involves sampling responses from strong language models, predominantly focusing on expanding the size of the dataset and improving the overall quality of the responses (Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024b; Chen et al., 2024). However, recent research suggests that there is more complex dynamics involved (Xu et al., 2024c). A plateau effect in synthetic data scaling, where performance either stagnates or even declines as the size of the synthetic data increases beyond a certain point, has been widely observed. This phenomenon arises due to issues such as diminishing diversity (Padmakumar & He, 2024; Guo et al., 2023) and distortion in the data distribution (LeBrun et al., 2021), which ultimately undermine the base model’s performance and robustness (Alemohammad et al., 2024; Gerstgrasser et al., 2024; Shumailov et al., 2023; Dohmatob et al., 2024; Hataya et al., 2023; Martínez et al., 2023a, b; Bohacek & Farid, 2023; Briesch et al., 2023).
@@
Thus, effective SFT requires more than scaling up the data; it often needs “tailoring” the data to the unique characteristics of the target model. Existing works focus on enhancing the model’s existing knowledge and capabilities (Du et al., 2023) and optimizing the curriculum progression for instruction tuning (Zhao et al., 2024b; Lee et al., 2024; Feng et al., 2023; Setlur et al., 2024). They typically find questions the model should best learn, instead of what answers the model should imitate.
Meanwhile, tailoring the responses to the target model has been a crucial ingredient for the success of later phases of LLM development, particularly through on-policy preference learning (Tajwar et al., 2024; Zhang et al., 2024c, a; Miao et al., 2024; Gulcehre et al., 2023; Azar et al., 2023; Tang et al., 2024a; Zhuang et al., 2023), and on-policy/online reinforcement learning (Guo et al., 2024b; Liu et al., 2024d; Zhou et al., 2024c).
@@
Inspired by these insights, we hypothesize that SFT can similarly benefit from aligning data with the model, the core idea behind GRAPE. For each instruction, GRAPE gathers and selects response(s) from various sources that are closest to the target model’s pretrained distribution. This is achieved by calculating the probability of each response using the target model and selects the one with the highest length-normalized probability(§3).After obtaining these more “in-distribution” responses, GRAPE proceeds with standard SFT without any modification to the training.
Unlike existing datasets that usually contains one-size-fits-all responses for each instruction without customization (Yu et al., 2024; Yuan et al., 2024b; Lian et al., 2023b; Teknium, 2023), GRAPE curates model-dependent SFT datasets that better matches the base model’s distribution, better mitigating the risks associated with distribution shift like spurious correlations (Zhou et al., 2024d) and catastrophic forgetting (Luo et al., 2025; Kotha et al., 2024), while posing minimum overhead of single forward pass over the candidate set. In return, GRAPE allows better downstream performance with reduced training compute.
@@
To GRAPE’s advantage, many existing datasets share overlapping instructions but contain different high-quality responses, e.g., the instructions in Flan (Longpre et al., 2023), GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and the post-training recipes that re-use SFT instructions for preference learning (Lambert et al., 2024; OLMo et al., 2025). Therefore, GRAPE can directly select, for each model, the best fit(s) among the off-the-shelf responses without having to produce new responses, which proves effective in our experiments (§ 5).
[…]
We then experiment with a more realistic setting for general-domain instruction-tuning, collecting responses from post-training data for Tulu3 and Olmo-v2. Again, GRAPE demonstrates its effectiveness by outperforming state-of-the-art data selection approaches by avg. 4.6% and a strong baseline trained on all these available data that is 4.5 times larger by up to 6.1%. Remarkably, GRAPE allows finetuning a Llama3.1-8B base model to exceed the performance of Tulu-8B-SFT using 1/3 data and half of the epochs.
Our results reveal that distributional alignment is a crucial and previously underappreciated dimension of effective instruction tuning. GRAPE introduces this perspective with a simple, scalable algorithm that consistently outperforms strong baselines with far less data and compute for the actual training.
@@ 2 Background and Motivation
2.1 Data Engineering for Instruction Tuning
Data is central to the success of effective instruction tuning,  (Xu et al., 2023; Xia et al., 2024; Chan et al., 2024), featuring both automated data synthesis (Xu et al., 2024a; Zeng et al., 2024; Yu et al., 2024; Wei et al., 2023) and selection (Xia et al., 2024; Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024e). Some selection approaches focus on high-quality data by leveraging LLMs (Chen et al., 2023a; Parkar et al., 2024; Li et al., 2024b) or employing principled metrics (Kang et al., 2024; Mekala et al., 2024; Xia et al., 2024), while others, such as  Yang et al. (2024b); Das & Khetan (2023), aim to identify diversity-optimized subsets for greater efficiency. An emerging trend is the customization of training data based on the characteristics of the base models. (Li et al., 2024c; Du et al., 2023; Li et al., 2024a). These methods typically reweight or filter instructions. However, a key overlooked aspect is response selection—specifically, choosing responses that align with the model’s pretrained distribution, which may be critical for preserving useful behaviors and ensuring effective fine-tuning.
Figure 1: Average performance curve of Llama3.1-8B on Tulu-3-Olmo-2 combo of GRAPE against random selection
2.2 Toward Distribution-Aligned Supervised Fine-Tuning
@@
The investigation of this work into the distribution match between the pre-trained LM and supervised fine-tuning (SFT) data is inspired by recent findings on policy optimization for LM alignment with RL (Ouyang et al., 2022) and preference learning (Rafailov et al., 2023; Ethayarajh et al., 2024). While the importance of matching training data distribution with the policy has been well noted in both traditional RL (Shi et al., 2023; Fujimoto et al., 2018; Kumar et al., 2019; Peng et al., 2019; Wang et al., 2021; Arora & Goyal, 2023; Jiang & Li, 2016; Tang & Abbeel, 2010) and LM settings (Xiong et al., 2024), preference learning algorithms like DPO (Rafailov et al., 2023), IPO (Azar et al., 2023) and KTO (Ethayarajh et al., 2024) first emerged as off-policy algorithms. However, subsequent research has highlighted the performance gap between on-policy and off-policy training due to distribution shifts  (xu2024dposuperiorppo; Tang et al., 2024b) and proposed various mitigation strategies  (Zhuang et al., 2023; Zhou et al., 2024c; Zhang et al., 2024a; Xiong et al., 2024; Guo et al., 2024b), showing that training models on data more closely aligned with their policy distribution can significantly improve performance, while failing to do so can yield sub-optimal policies or those that are harder to generalize. Works like SPIN (Yuan et al., 2024a) echo the intuition by gradually improving policy through self-play to mitigate distribution shift.
Figure 2: Training loss curve of Llama3.1-8B on Tulu-3-Olmo-2. Best (highest-probability, selected by GRAPE), Random, and Worst (lowest-probability) responses show a clear loss ordering: Best << Random << Worst.
@@
The base distribution of pre-trained language models—shaped by extensive training on vast and diverse datasets—is inherently robust and generalizable (Brown et al., 2020; Saunshi et al., 2021). Therefore, during supervised fine-tuning phase, the pre-trained distribution should be carefully preserved (Kumar et al., 2022; Cohen-Wang et al., 2024; He et al., 2023; Yang et al., 2024d; Ding et al., 2023), to best retain the knowledge and capabilities that emerge during pre-training (Zhou et al., 2023). If the proximity between the pre-trained distribution and the fine-tuning data is not maintained, the limited number of training examples available during SFT, compared to the vast scale of pre-training data, can increase the risk of distribution distortion. This misalignment can lead to issues such as catastrophic forgetting (Aghajanyan et al., 2020; Yang et al., 2024d) and the emergence of spurious correlations (Feldman, 2021).
@@
We build on principles of on-policy alignment techniques with key distinctions tailored for SFT. Yet, given that SFT represents an earlier stage of post-training than RL, sampling responses from the base model itself alone can lead to model collapse, as noted in (Shumailov et al., 2024). Prior studies have documented similar risks like instability, bias reinforcement, knowledge stagnation, and overfitting (Herel & Mikolov, 2024; Mobahi et al., 2020; Allen-Zhu & Li, 2023; Ghosh et al., 2024; Dong et al., 2025; Zhang et al., 2024d). To address this, we advocate for an approach that stays more in-distribution while delivering effective supervision to the base model. To this end, we propose to gather and select responses from various sources, and select one that is closest to the target model’s pretrained distribution, which we name GRAPE.
@@ 3 Methodology
We introduce GRAPE, a surprisingly simple yet effective methodology to enhance supervised fine-tuning (SFT) by customizing the training data for the base model. The key idea is to find a response, among a candidate pool, for each instruction $x_{i}$ that aligns closely with the base model’s pretrained distribution $\pi_{\theta_{0}}$.
@@
Response Collection (§3.1) Collect a pool of high-quality candidate responses from various sources.
@@ 3.1 Collecting Responses from Existing Resources
For instruction-tuning of language models, high-quality instructions are more difficult to collect than responses (Xu et al., 2024b; Liu et al., 2024a). Therefore, it is a common practice to reuse existing instruction-tuning prompts while generating diverse responses using various methods tailored to specific requirements. For instance, instructions from Flan (Longpre et al., 2023), OpenOrca (Lian et al., 2023a), ShareGPT (Team, 2023), and the training splits of GSM-8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and CodeContests (Li et al., 2022) are frequently reused in datasets like Olmo (OLMo et al., 2025), Tulu (Lambert et al., 2024), OpenHermes (Teknium, 2023), OpenOrca (Lian et al., 2023a), MetaMath (Yu et al., 2024), MathInstruct (Yue et al., 2023), UltraFeedback (Cui et al., 2024), and UltraInteract (Yuan et al., 2024b), whether for SFT or preference learning. The solutions are generated using different models or follow varying styles depending on the specific needs. This naturally leads to a situation where a single instruction with multiple responses becomes a readily available resource. GRAPE therefore leverages these pre-existing response candidates to tailor training dataset that better aligns with the base model’s distribution. When such resources are unavailable or insufficient, practitioners can generate new responses and apply GRAPE.
For each instruction, we collect multiple responses from various datasets. These responses form a candidate set associated with the instruction.
@@ 3.2 Customize Dataset For Models
[…]
Since GRAPE only involves forward-pass log-probability computation (no gradients or optimization), it is highly efficient and simple to integrate into any SFT pipeline with minimal overhead compared with model-based data selection approaches (Xia et al., 2024; Yang et al., 2024b; Liu et al., 2024b; Zhao et al., 2021; Zhang et al., 2024b; Pan et al., 2024). Additionally, it is important to distinguish GRAPE from the other perplexity-based data selection and curriculum planning methods (Wu et al., 2024; Li et al., 2024b; Liu et al., 2024c). Existing approaches focus on selecting instructions by using perplexity as a difficulty measure, which differs from GRAPE that uses probability to select for each instruction in a fixed instruction set, responses that better matches with the base model’s distribution. Our experiments in §5 demonstrate that low-probability responses with fixed set of instructions are detrimental to performance, further emphasizing the fundamental difference in the two processes.
4 Controlled Experiments Show the Benefits of Distributional Alignment
We conduct controlled experiments using UltraInteract-SFT to test whether selecting responses aligned with a base model’s distribution improves fine-tuning outcomes more than relying on stronger generators or scaling data size. This setup—focused on verifiable tasks in coding, logic, and math—isolates the effect of distribution matching. Results show that GRAPE-selected responses consistently outperform alternatives, validating distributional alignment as a key supervision signal. These findings motivate our broader evaluations in §5.
@@ 4.1 Experimental Setup
Training Data Curation
In this controlled experiment, we focus on chain-of-thought reasoning (Wei et al., 2022; Wang et al., 2024a; Luo et al., 2024; Cobbe et al., 2021; Li et al., 2023; Lightman et al., 2023). Different models may follow different reasoning paths to solve a problem, while their final solutions can be easily verified.
We use UltraInteract-SFT (Yuan et al., 2024b), which contains approximately $80,800$ unique instructions covering coding, math (chain-of-thought and program-aided) and logic reasoning domains , where each instruction is paired with varying numbers (avg. 3.5/instruction) of different responses to contain a total >280,000>280,000 training examples. The responses in the dataset are strictly in step-wise format. For each instruction, we construct a response pool consisting of both (i) original UltraInteract-SFT responses and (ii) additional responses generated from a diverse set of LLMs. GRAPE is then applied to select the most in-distribution response per instruction to this enlarged candidate pool and ensure the number of responses matches the original UltraInteract-SFT dataset for fair comparisons.
We collect responses from a diverse set of models of various sizes across model families, including Mixtral-7x7B-Instruct (Jiang et al., 2024), Codestral-22B (MistralAI, 2024a), Mistral-Small (MistralAI, 2024b), Llama-3.1-70B-Instruct and Llama-3.1-405B-Instruct (Dubey et al., 2024), and Qwen2.5-72B-Chat (Yang et al., 2024a), resulting in approximately 10x additional responses per instruction. The responses are then filtered based on the answers to ensure their validity following Yuan et al. (2024b).
Base Models
To demonstrate the generalizability of GRAPE, we evaluate its performance across multiple LLMs, including Llama-3.1-8B and Llama-3.2-3B from Llama-3 (Grattafiori et al., 2024) family, Mistral-7B (Jiang et al., 2023) and Qwen2.5-7B (Hui et al., 2024). We ensure that all training configurations (GRAPE, baselines) use the same number of instructions and responses per instruction as original UltraInteract-SFT, unless otherwise stated.
Evaluation
@@
Baselines
[…]
@@
Qwen2.5-7B GRAPE 77.4 48.9 70.7 56.4 45.3 67.7 66.3 37.4 40.1 56.7 -
Table 1: Result of synthetic experiment on UltraInteract-SFT. The last column, Abs. $\Delta$ is GRAPE’s absolute improvement on average performance over that row.